Original observed population:

No stripes: 2

One stripe:5

Many stripes: 23

Assume population is under selection, you are then left with a population of breeders.

Selection coefficient(s) acting on each phenotype.

0 bands = 0.5

1 band = 0.3

Many bands = 0

What are the phenotypic and genotypic frequencies of the breeders?
Answer #1

Selection is generally defined on the basis of fitness (w), where fitness is given by
w= 1-s where s is selection coefficient.
So, for
0 bands, w = 1-0.5 = 0.5
1 band w = 1-0.3 = 0.7
Many bands w = 1-0 = 1.0

Here the fitness is relative fitness which signifies the percentage of breeders in a population surviving.
So, this implies for
0 band probability of survival is 0.5 x 100 = 50%
1 band probability of survival is 0.7 x 100 = 70%
many bands probability of survival is 1 x 100 = 100%

So the frequency of breeders would be
0 band = 2 x 50/100 = 1
1 band = 5 x 70/100 = 3.5
Many bands = 23 x 100/100 = 23